OpenStudy (anonymous):

Find the inverse function for f(t) = 50e^0.25t

OpenStudy (anonymous):

I got ln(x/50)/0.25 but its wrong? what did i do wrong?

OpenStudy (anonymous):

it is non invert able

OpenStudy (anonymous):

mention range and domain to truncate it in its domain

OpenStudy (anonymous):

but there has to be since non invertable is not an answer

OpenStudy (unklerhaukus):

\[f(t) = 50e^{0.25t}\qquad f>0\] \[\frac{f(t)}{50}=e^{0.25t}\] \[\ln\frac{f(t)}{50}=0.25t\] \[4\ln\frac{f(t)}{50}=t\] \[t(f)=4\ln\frac{f}{50}\qquad t>0\]